# Cross-lingual Offensive Language Detection: A Systematic Review of Datasets, Transfer Approaches and Challenges

### v1

@@ 6.3.2. Hybrid Transfer Strategies
This strategy, often referred to as “learning to learn”, has become a popular few-shot learning technique in NLP. It involves training models to learn the optimal initialisation of parameters, allowing them to be fine-tuned with a small amount of data from the target language. This technique finds its application within the joint learning and cascade learning transfer scenarios. It enables rapid adaptation to new unseen languages, making it especially useful in scenarios where labelled data is limited. The main meta learning methods include optimisation-based (Finn et al., 2017) and metric-based techniques (Koch et al., 2015). Two papers in our review employ optimisation-based meta learning framework. Montariol et al. (2022) study meta learning for the problem of few-shot hate speech detection in low-resource languages. They propose a cross-lingual meta learning-based approach based on optimisation-based Model-Agnostic Meta-Learning (MAML) and Proto-MAML models, fine-tuning the base learner XLM-R with parallel few-shot datasets in different target languages. Awal et al. (2023) propose HateMAML, a model-agnostic meta-learning-based framework that uses a semi-supervised self-refinement strategy to fine-tune a better PLM for unseen data in target language, showing effective performance for hate speech detection in low-resource languages.
Ensemble Learning
@@
It leverages multiple models or “learners” to collaboratively make decisions. Instead of relying on the output of a single model, ensemble learning combines all predictions from multiple models to produce a final prediction based on majority voting (Paul et al., 2023; Markov et al., 2021). It takes advantage of the different strengths of each model, mitigating the weaknesses and biases of any single model and avoiding overfitting, improving robustness and reliability over a single classifier. This is particularly beneficial in cross-lingual settings, where language biases can pose challenges to individual models. In cross-lingual offensive language detection, ensemble learning is mainly applied in the zero-shot, joint learning, and cascade learning transfer scenarios. Deep ensemble models have emerged as the predominant choice in recent studies (Bigoulaeva et al., 2021, 2022, 2023; Paul et al., 2023), while the integration of PLMs within ensemble learning frameworks is also common (Burtenshaw and Kestemont, 2021; Ryzhova et al., 2022; Markov et al., 2021). Each model in the ensemble can be trained on different languages or linguistic features, allowing the ensemble model to capture diverse linguistic patterns and reduce test errors for target languages (Ryzhova et al., 2022).
@@ 6.4. Summary of Cross-lingual Approaches
For transfer scenarios, zero-shot transfer does not use any target data, while joint and cascade learning do. With zero-shot learning, we always transfer the whole model, while with joint and cascade learning we might transfer only a subset of parameters. Some overlaps of transfer strategies might exist between each transfer scenario because multiple strategies can be conducted based on more than one transfer scenario.
Model distributions. We provide an overview of diverse models and their frequency used in recent cross-lingual studies, which reveals a clear trend towards more advanced architectures. Traditional machine learning-based models have been employed 24 times, while deep learning-based models have seen a slightly higher usage (38 times) due to their complex structure and learning capacity from a large amount of data. Transformer-based models have been overwhelmingly preferred (99 times), highlighting their growing dominance in NLP. Large Language Models have also recently been investigated in this area due to their excellent emergent ability.
Evaluation metrics. To evaluate model performance in cross-lingual offensive language detection, macro F1 is the most used metric given class imbalance in datasets, while accuracy is also popular. Some studies also consider precision, recall, weighted F1 score for a detailed assessment. Eronen et al. (2022) propose a new linguistic similarity metric based on the World Atlas of Language Structures (WALS) (Dryer and Haspelmath, 2013) to select optimal transfer languages for automatic hate speech detection.
@@
Language pairs. We present the frequency of language pairs considered in cross-lingual studies in Figure 7. English prevails as the source language, while the most frequent target languages are Spanish, German, and Italian. These languages share typological similarities with English, which likely facilitates their frequent pairing in cross-lingual hate speech detection. This emphasises the impact of data availability and language similarity on the training and application of cross-lingual methods, while also revealing the limitation of relevant cross-lingual research resulting from data scarcity. Many studies have also explored cross-lingual learning by transferring English to Arabic, Hindi and French, as these languages have a relatively large amount of datasets available (see Figure 7). Additionally, there are studies working on the transfer between non-English languages, among which Danish, Indonesian, Portuguese, and Turkish are frequent.
@@ 7.1.1. Diverse Linguistic Structures
Language diversity encompasses various aspects from basic grammatical rules to complex nuances unique to each language. Since languages within a certain language family may exhibit similarities, cross-lingual transfer works better in more closely related languages and poses challenges for more dissimilar languages (R”̈ottger et al., 2022b). However, this does not necessarily imply a consistency in linguistic features (Eronen et al., 2022). Furthermore, the presence of dialectal forms within a single language amplifies its complexity. For example, the Arabic language is spoken by a wide range of countries across Asia, with complexities arising due to its multiple dialects (Husain and Uzuner, 2022).
@@ 7.1.2. Code Mixing
Non-English social media content is often characterised by code-mixing, where users blend languages, use transliterations, or incorporate multiple scripts within a single sentence or conversation, such as English-Hindi (Ibrohim and Budi, 2019a; Hande et al., 2021b) and English-Chinese (Jiang and Zubiaga, 2023). The English-Hindi combinations (also called Hinglish) appear in many cases, because the unique policies on regulating such speech in these regions add layers of complexity and make detection more challenging (Kapoor et al., 2019).
@@ 7.1.3. Cultural Variations
Culture is multifaceted and complex. The perception of offensive language can differ significantly across languages and cultures (Nozza, 2021). What is considered non-offensive in one culture or language might be misinterpreted as signals of offense in another, leading to detection discrepancies. Even within a single language, there can be cultural diversity. For example, categorising English as a representative of “western cultural background” might ignore the differences between American and British cultures (Zhou et al., 2023). Such cultural variations can hinder the effective transferability of language models (R”̈ottger et al., 2022b). Moreover, hateful expressions often employ figurative language, rhetorical figures and idioms, which are also language-dependent. This requires models to interpret underlying intent rather than relying solely on literal meanings (Pamungkas et al., 2021a).
@@ 7.2. Dataset-Related Challenges
7.2.1. Limited Labelled Datasets
Low-resource languages often do not have the attention that major languages receive, leading to limited data availability (Burtenshaw and Kestemont, 2021) and limited studies (see Figures 3 and 7). Additionally, creating a high-quality labelled dataset requires extensive manual effort, which can potentially lead to high annotation and time costs as well as be harmful to annotators (R”̈ottger et al., 2022a). Annotating offensive language is even more challenging due to the need for cultural and linguistic knowledge. Ethical and privacy concerns also arise when dealing with real-world data from social media platforms or other online sources (Jiang et al., 2022), because there are potential risks of exposing sensitive information or inadvertently promoting offensive language when studying it or trying to combat it.
@@ 7.2.2. Inconsistent Definition of Offensive Language
The diversity of hate speech encompassing from misogyny and racism to other forms of discrimination often leads to discrepancies in research (Nozza, 2021). This has a bigger impact on low-resource languages (Bigoulaeva et al., 2021, 2022, 2023), and introduce ambiguities that challenge the task, especially in multilingual scenarios (Vitiugin et al., 2021). The lack of definition consistency across datasets in turn limits cross-lingual research, because many datasets in either source or target languages may be incompatible (Bigoulaeva et al., 2022).
7.2.3. Inconsistent Annotations
Offensive language datasets can exhibit systematic gaps and biases in their annotations (R”̈ottger et al., 2022b). Different labelled datasets often utilise distinct annotation frameworks or strategies (Fortuna and Nunes, 2018). The subjective nature of offensive language leads to tentatively disputed annotations among human annotators, where hateful content could be interpreted differently depending on annotators’ understanding (Das et al., 2022). Annotating these contents is easily influenced by individual sociodemographic and cultural backgrounds, often resulting in low inter-annotator agreement (Markov et al., 2021). Additionally, there is an “annotation’s dilemma” (Aluru et al., 2021), whereby ambiguous instances might be annotated incorrectly by annotators but predicted accurately by the model.
7.2.4. Dataset Imbalance
@@
Given that the majority of social media content is non-hateful, labels are typically skewed towards the non-hate label (Song et al., 2021; Bigoulaeva et al., 2022, 2023). This skewness can lead to training issues, especially when working with small training corpora. In addition, when considering multilingual or merged datasets, inter-language imbalance is another challenge (Deshpande et al., 2022). Such datasets often display significant differences in the number of examples available for each language. This imbalance can bias models towards the semantic tendencies of languages that are overrepresented.
7.2.5. Dataset Bias
@@
Biases such as topic, authorship, political affiliation, social media platform, and collection period can be one of the major issues and lead to potential generalisation challenges for cross-lingual models. Among them, topic bias has been taken seriously in many works (R”̈ottger et al., 2022a; Arango et al., 2022; Vitiugin et al., 2021; R”̈ottger et al., 2022b). While some datasets include general offensive language or hate speech, others might concentrate on specific issues such as immigrants, misogyny, politics, or religion. Such topic-specific biases across datasets can harm model performance in cross-lingual classification (Pamungkas et al., 2021a; Stappen et al., 2020). Besides, the temporal aspect of data collection, affected by events in different periods, can further introduce bias, leading to datasets with varied topical foci (Stappen et al., 2020).
7.2.6. Data Source Obsolescence
@@
As online language rapidly transforms, especially influenced by moderation and real-world events, datasets can quickly become obsolete (Vitiugin et al., 2021; Alkhalifa et al., 2023). This obsolescence is further exacerbated by the emergence of new slang, metaphors, and colloquialisms that vary across languages and regions. For instance, terms that normally have neutral meanings (such as donkey) may be used offensively in sexist text (Jiang and Zubiaga, 2023). The constrained and informal style of tweets, which are more like oral expressions than written language, complicates the preprocessing steps and leads to erroneous predictions (Pamungkas et al., 2020; Aluru et al., 2021). Emojis, associated with various forms of online harassment, further pose unique challenges (Corazza et al., 2020). Moreover, the domain and target of hate speech can shift significantly over time (Florio et al., 2020). Real-world events, from local incidents to global crises like the COVID-19 pandemic, can give rise to new domains and terms of hate speech (Montariol et al., 2022). Then, it becomes increasingly challenging to create and continually update datasets customised for every possible language and domain, leading to frequent low-resource issues.
@@ 7.3.1. Limited Ability of Multilingual Pre-Trained Language Models
Cross-lingual models, especially multilingual PLMs, often face challenges related to their generalisation abilities (Hahn et al., 2021; Pelicon et al., 2021b). The generalisability of transformer-based multilingual PLMs can be inconsistent, especially for typologically diverse languages, because they might be pre-trained with a highly focused set of languages or some languages with insufficient training samples (Deshpande et al., 2022). This can lead to bias and instability in model performance, depending on the model architectures or topical focus in datasets (Mozafari et al., 2022; Pamungkas et al., 2021a). Model overfitting across target languages is another important issue. Although multilingual models are less prone to overfitting on dataset-specific features than monolingual models, they will achieve poorer performance than monolingual ones in higher-resource settings, and they may require very different calibrations and adaptations across languages (R”̈ottger et al., 2022a, b).

### v2

@@ 6.3.2. Hybrid Transfer Strategies
Meta Learning. This strategy, often referred to as “learning to learn”, has become a popular few-shot learning technique in NLP. It involves training models to learn the optimal initialisation of parameters, allowing them to be fine-tuned with a small amount of data from the target language. This technique finds its application within the joint learning and cascade learning transfer scenarios. It enables rapid adaptation to new unseen languages, making it especially useful in scenarios where labelled data is limited. The main meta learning methods include optimisation-based (Finn et al., 2017) and metric-based techniques (Koch et al., 2015). Two papers in our review employ optimisation-based meta learning framework. Montariol et al. (2022) study meta learning for the problem of few-shot HS detection in low-resource languages. They propose a cross-lingual meta learning-based approach based on optimisation-based Model-Agnostic Meta-Learning (MAML) and Proto-MAML models, fine-tuning the base learner XLM-R with parallel few-shot datasets in different target languages. Awal et al. (2023) propose HateMAML, a model-agnostic meta-learning-based framework that uses a semi-supervised self-refinement strategy to fine-tune a better PLM for unseen data in target language, showing effective performance for HS detection in low-resource languages.
@@
Ensemble Learning. It leverages multiple models or “learners” to collaboratively make decisions. Instead of relying on the output of a single model, ensemble learning combines all predictions from multiple models to produce a final prediction based on majority voting (Paul et al., 2023; Markov et al., 2021). It takes advantage of the different strengths of each model, mitigating the weaknesses and biases of any single model and avoiding overfitting, improving robustness and reliability over a single classifier. This is particularly beneficial in cross-lingual settings, where language biases can pose challenges to individual models. In cross-lingual offensive language detection, ensemble learning is mainly applied in the zero-shot, joint learning, and cascade learning transfer scenarios. Deep ensemble models have emerged as the predominant choice in recent studies (Bigoulaeva et al., 2021, 2022, 2023; Paul et al., 2023), while the integration of PLMs within ensemble learning frameworks is also common (Burtenshaw and Kestemont, 2021; Ryzhova et al., 2022; Markov et al., 2021). Each model in the ensemble can be trained on different languages or linguistic features, allowing the ensemble model to capture diverse linguistic patterns and reduce test errors for target languages (Ryzhova et al., 2022).
@@ 6.4. Summary of Cross-lingual Approaches
A review of the surveyed work reveals several key developments in cross-lingual offensive language detection. First, there has been a rise of multilingual PLMs, beginning with general-purpose PLMs such as mBERT and XLM-R, which quickly replaced feature- or translation-based methods as the dominant approach (Devlin et al., 2019; Conneau et al., 2020). Meanwhile, there is a move from general-purpose PLMs to more language- and culture-specific PLMs, such as MuRIL for Indic languages (Das et al., 2022), AraBERT for Arabic dialects (Husain and Uzuner, 2022), and XLM-T trained on social media (Montariol et al., 2022). More recently, a new trend has emerged with the rise of large language models (LLMs), capable of zero- or few-shot multilingual tasks via prompting (Brown et al., 2020; OpenAI, 2023; Plaza-del arco et al., 2023). See future challenges and directions in PLMs and LLMs in §7.4. Second, we observe a shift from feature-based to parameter-based transfer. Earlier approaches relied on CLWEs, lexicons, or projection methods (Pamungkas and Patti, 2019; Arango et al., 2021). With the rise of multilingual PLMs, parameter-based fine-tuning has become the core transfer mechanism, while handcrafted features now play a complementary role. Moreover, there is a growing focus on low-resource languages and dialects. While early research centred on English and a few European languages, more recent work increasingly covers Indic (Das et al., 2022), Arabic (Husain and Uzuner, 2022), and Asian languages (Jiang et al., 2022; Jeong et al., 2022). However, many languages still lack dedicated datasets and evaluation benchmarks.
Model distributions. After reviewing diverse models and their frequency of use in recent studies, we can see traditional machine learning-based models have been employed 24 times, while deep learning-based models have seen a slightly higher usage (38 times) due to their complex structure and learning capacity from a large amount of data. Transformer-based models have been overwhelmingly preferred (99 times), highlighting their growing dominance in NLP. Large Language Models have also recently been investigated in this area due to their excellent emergent ability.
Evaluation metrics. Macro F1 is the most used metric given class imbalance in datasets, while accuracy is also popular. Some studies also consider precision, recall, and weighted F1 score. Eronen et al. (2022) propose a new linguistic similarity metric based on the World Atlas of Language Structures (WALS) (Dryer and Haspelmath, 2013) to select optimal transfer languages for automatic HS detection.
@@
Language pairs. We present the frequency of language pairs in Figure 7. English prevails as the source language, while the most frequent target languages are Spanish, German, and Italian. These languages share typological similarities with English, which likely facilitates their frequent pairing in cross-lingual tasks. This emphasises the impact of data availability and language similarity on the training and application of cross-lingual methods, and also reveals the limitations resulting from data scarcity. Many studies have also explored cross-lingual learning by transferring English to Arabic, Hindi and French, as these languages have a relatively large amount of datasets available (see Figure 3). Additionally, there are studies working on the transfer between non-English languages, among which Danish, Indonesian, Portuguese, and Turkish are frequent.
Comparative strengths and trade-offs. While each transfer level offers viable solutions, their effectiveness can vary depending on the task setting and language pair characteristics. For example, instance transfer is often most effective when high-quality parallel corpora or reliable translation tools are available, enabling direct label projection or text translation. It tends to work well for typologically similar languages or when cultural and topical overlap is high. Feature transfer is particularly useful when parallel data is scarce but high-quality multilingual representations exist; it can capture shared semantic spaces while retaining language-specific nuances. This approach is generally more robust to moderate typological distance, but may struggle when the target language is under-represented in pre-training corpora. Parameter transfer offers the best flexibility across settings, but its success depends heavily on the pre-trained model coverage of the target language and the domain similarity between training and testing data. Parameter sharing in multilingual PLMs often generalises best within the same language family, while cross-family transfer may require additional adaptation (e.g., domain-specific fine-tuning or language-specific modules). In practice, selecting an appropriate transfer strategy requires balancing data availability, typological similarity between languages, and computational resources. Hybrid approaches that combine strategies can often yield the most reliable performance across diverse language families, such as using instance transfer for initial alignment and parameter transfer for fine-tuning.
@@ 7.1. Language-Related Challenges
7.1.1. Low-resource and Underrepresented Languages
A central challenge in cross-lingual offensive language detection is the limited availability of resources for many non-mainstream languages. As illustrated in Figure 7, the majority of existing studies focus on English as the source language and transfer primarily to typologically similar European languages such as Spanish, German, Italian, and French. In contrast, languages from other families (particularly African, Indigenous, and smaller Asian languages) remain severely underrepresented in both datasets and modelling efforts. This imbalance reflects broader trends in multilingual PLMs, where training corpora disproportionately favour high-resource languages (Nozza, 2021; Röttger et al., 2022a). As a result, CLTL systems often exhibit strong performance for well-represented languages but much weaker transferability for low-resource or distant languages (Deshpande et al., 2022; Eronen et al., 2022). The scarcity of labelled data, the uneven distribution of languages in PLMs, and the difficulty of transferring knowledge across typologically distant families represent major obstacles.
@@ 7.1.2. Diverse Linguistic Structures
Language diversity encompasses various aspects from basic grammatical rules to complex nuances unique to each language. Linguistic distance captures how typologically similar or different two languages are. Transfer tends to be more successful across languages within a certain family that shares vocabulary, morphology, or syntactic patterns (e.g., English–German, Spanish–Italian), while performance often degrades for distant language pairs (e.g., English–Swahili, English–Chinese) (Deshpande et al., 2022; Eronen et al., 2022). Figure 7 also demonstrates that cross-lingual experiments are concentrated within Indo-European languages, while distant language pairs are far less represented. This indicates that challenges for CLTL are not only the scarcity of data in low-resource languages, but also the uneven ability of models to generalise across varying degrees of linguistic distance. Furthermore, the presence of dialectal forms within a single language amplifies its complexity. For example, the Arabic language is spoken by a wide range of countries across Asia, with complexities arising due to its multiple dialects (Husain and Uzuner, 2022).
@@ 7.1.3. Code Mixing
Non-English social media content is often characterised by code-mixing, where users blend languages, use transliterations, or incorporate multiple scripts within a single sentence or conversation, such as English-Hindi (Ibrohim and Budi, 2019a; Hande et al., 2021b) and English-Chinese (Jiang and Zubiaga, 2023). The English-Hindi (or Hinglish) combination appears in many cases, because the unique policies on regulating such speech in these regions add layers of complexity and make detection more challenging (Kapoor et al., 2019).
@@ 7.1.4. Cultural Variations
Culture is multifaceted and complex. The perception of offensive language can differ significantly across languages and cultures (Nozza, 2021). What is considered non-offensive in one culture or language might be misinterpreted as signals of offense in another, leading to detection discrepancies. Even within a single language, there can be cultural diversity. For example, categorising English as a representative of “western cultural background” might ignore the differences between American and British cultures (Zhou et al., 2023). Such cultural variations can hinder the effective transferability of language models (Röttger et al., 2022b). Moreover, hateful expressions often employ figurative language, rhetorical figures and idioms, which are also language-dependent. This requires models to interpret underlying intent rather than relying solely on literal meanings (Pamungkas et al., 2021a).
7.1.5. Explicit vs. Implicit Offensive Language
Offensive language can be expressed explicitly or implicitly. Explicit abuse, such as direct insults or slurs, often transfers more easily across languages because it relies on surface lexical markers. In contrast, implicit abuse (expressed through sarcasm, stereotypes, or code-mixed language) is far more context-dependent and culturally situated. As a result, implicit forms are much harder to transfer from high-resource to low-resource settings, where contextual cues may differ significantly (Hartvigsen et al., 2022).
@@ 7.2.1. Inconsistent Definition of Offensive Language
The diversity of HS encompassing from misogyny and racism to other forms of discrimination often leads to discrepancies in research (Nozza, 2021). This has a bigger impact on low-resource languages (Bigoulaeva et al., 2021, 2022, 2023), and introduces ambiguities that challenge the task, especially in multilingual scenarios (Vitiugin et al., 2021). The lack of definition consistency across datasets in turn limits cross-lingual research, because many datasets in either source or target languages may be incompatible (Bigoulaeva et al., 2022).
7.2.2. Data Scarcity and Quality
Limited Labelled Datasets. Low-resource languages often do not have the attention that major languages receive, leading to limited data availability (Burtenshaw and Kestemont, 2021) and limited studies (see Figures 3 and 7). Additionally, creating a high-quality labelled dataset requires extensive manual effort, which can potentially lead to high annotation and time costs as well as be harmful to annotators (Röttger et al., 2022a). Annotating offensive language is even more challenging due to the need for cultural and linguistic knowledge. Ethical and privacy concerns also arise when dealing with real-world data from social media platforms or other online sources (Jiang et al., 2022), because there are potential risks of exposing sensitive information or inadvertently promoting offensive language when studying it or trying to combat it.
@@ 7.2.2. Data Scarcity and Quality
Dataset Imbalance. Given that the majority of social media content is non-hateful, labels are typically skewed towards the non-hate label (Song et al., 2021; Bigoulaeva et al., 2022, 2023). This skewness can lead to training issues, especially when working with small training corpora. In addition, when considering multilingual or merged datasets, inter-language imbalance is another challenge (Deshpande et al., 2022). Such datasets often display significant differences in the number of examples available for each language. This imbalance can bias models towards the semantic tendencies of languages that are overrepresented.
@@
Dataset Bias. Biases such as topic, authorship, political affiliation, social media platform, and collection period can be one of the major issues and lead to potential generalisation challenges for cross-lingual models. Among them, topic bias has been taken seriously in many works (Röttger et al., 2022a; Arango et al., 2022; Vitiugin et al., 2021; Röttger et al., 2022b). While some datasets include general offensive language or HS, others might concentrate on specific issues such as immigrants, misogyny, politics, or religion. Such topic-specific biases across datasets can harm model performance in cross-lingual classification (Pamungkas et al., 2021a; Stappen et al., 2020). Besides, the temporal aspect of data collection, affected by events in different periods, can further introduce bias, leading to datasets with varied topical foci (Stappen et al., 2020).
@@
Data Source Obsolescence. As online language rapidly transforms, especially influenced by moderation and real-world events, datasets can quickly become obsolete (Vitiugin et al., 2021; Alkhalifa et al., 2023). This obsolescence is further exacerbated by the emergence of new slang, metaphors, and colloquialisms that vary across languages and regions. For instance, terms that normally have neutral meanings (such as donkey) may be used offensively in sexist text (Jiang and Zubiaga, 2023). The constrained and informal style of tweets, which are more like oral expressions than written language, complicates the preprocessing steps and leads to erroneous predictions (Pamungkas et al., 2020; Aluru et al., 2021). Emojis, associated with various forms of online harassment, further pose unique challenges (Corazza et al., 2020). Moreover, the domain and target of HS can shift significantly over time (Florio et al., 2020). Real-world events, from local incidents to global crises like the COVID-19 pandemic, can give rise to new domains and terms of HS (Montariol et al., 2022). Then, it becomes increasingly challenging to create and continually update datasets customised for every possible language and domain, leading to frequent low-resource issues.
7.2.3. Annotation Bias
Offensive language datasets can exhibit systematic gaps and biases in their annotations (Röttger et al., 2022b). First, different labelled datasets often utilise distinct annotation frameworks or strategies (Fortuna and Nunes, 2018). Second, the subjective nature of offensive language leads to tentatively disputed annotations among human annotators, where hateful content could be interpreted differently depending on annotators’ understanding (Das et al., 2022). The annotation of these contents is easily influenced by individual sociodemographic and cultural backgrounds, often resulting in low inter-annotator agreement (Markov et al., 2021). Additionally, there is an “annotation dilemma” (Aluru et al., 2021), whereby ambiguous instances might be annotated incorrectly by annotators but predicted accurately by the model.
@@ 7.3.1. Limited Ability of Multilingual Pre-Trained Language Models
Cross-lingual models, especially multilingual PLMs, often face challenges related to their generalisation abilities and model bias (Pamungkas et al., 2021a; Hahn et al., 2021; Pelicon et al., 2021b). The generalisability of transformer-based multilingual PLMs can be inconsistent, especially for typologically diverse languages, because they might be pre-trained with a highly focused set of languages or some languages with insufficient training samples (Deshpande et al., 2022). This can lead to bias and instability in model performance, depending on the model architectures, topical focus, linguistic distance, or cultural context in datasets (Mozafari et al., 2022; Pamungkas et al., 2021a). For example, PLMs may exhibit systematic bias when handling dialects, minority languages with various linguistic distances, and content towards marginalised communities. Since such varieties could be under-represented in pre-training corpora, PLMs tend towards standardised forms of dominant languages. This can result in degraded model performance, erasure of dialect-specific expressions, or misinterpretation of community-specific discourse. Model overfitting across target languages is another important issue. Although multilingual models are less prone to overfitting on dataset-specific features than monolingual models, they will achieve poorer performance than monolingual ones in higher-resource settings, and they may require very different calibrations and adaptations across languages (Röttger et al., 2022a, b).